# A positive meshless finite difference scheme for scalar conservation laws with adaptive artificial viscosity driven by fault detection

### v1

@@ 6.2 Example 2: Inviscid Burgers’ equation with smooth initial condition
In spite of the initial condition being smooth, a shock appears as time advances. Note that this problem is an adaptation of the example given in Section 7.1.1 of Hesthaven [2018], where the flux is chosen slightly differently as ${\bf F}(u)=u^{2}{\bf v}$. While no explicit exact solution is available, we consider as a reference a high-resolution solution obtained by a fifth order WENO scheme on the $400\times 400$ Cartesian grid, see the book Hesthaven [2018], Chapter 11 and its supplementary code BurgersWENO2D.m. For comparison with our method we choose a first order monotone scheme (Hesthaven [2018], Script 7.2 and code BurgersM2D.m) on the $200\times 200$ Cartesian grid with spacing $h=0.0025$. Contour lines of the reference solution and the solution by the first order monotone method at time $T=0.1$ are shown in Figure 6.4(a,b).
@@
The contour lines of the solutions at time $T=0.1$ obtained by our method are presented in Figure 6.4(c,d). In addition, we compare the cross sections of all four solutions for $x_{2}=0.375$ in Figure 6.5, and the errors with respect to the reference solution in Table 6.2. Note that in order to compute the cross sections and contour lines in the case of Halton nodes we interpolate the solution on the $200\times 200$ Cartesian grid by the piecewise linear polynomial on the Delaunay triangulation of the nodes. The results of the meshless scheme on Cartesian and Halton nodes are very similar, and significantly better than those of the monotone scheme. This is clear from the errors in the table and zoom-in plots in Figure 6.5(c-f). In particular, these results confirm remarkable robustness of our method with respect to the use of scattered points. Note that the contours in Figure 6.4(d) appear wobbling at the shocks, due to the irregular distribution of the Halton nodes in close proximity of the shock lines.
method $E_{1}$-error $E_{2}$-error monotone scheme (Hesthaven [2018], Script 7.2) 1.21e-02 4.38e-02 meshless scheme (Algorithm 3), Cartesian grid 6.08e-03 2.79e-02 meshless scheme (Algorithm 3), Halton points 5.47e-03 2.71e-02
Table 6.2: Example 2: Comparison of the errors (with respect to the reference solution) obtained by a standard monotone scheme on the Cartesian grid, and our meshless method on both Cartesian grid and Halton points ($T=0.1$).
@@
(c) Zoom no. 1.
(d) Zoom no. 2.
(e) Zoom no. 3.
(f) Zoom no. 4.
Figure 6.5: Example 2: Comparison of the cross sections at $x_{2}=0.375$ of the reference solution and the monotone scheme with the meshless scheme on (a) the Cartesian grid and (b) Halton points ($T=0.1$). Subfigures (c)-(f) show zoom-in plots of four critical portions of the cross sections near the shocks.
@@ 6.3 Example 3: Rotating wave
$u_{0}({\bf x})=\left\{\begin{array}[]{ll}3.25\pi,&\mbox{if }\|{\bf x}\|_{2}<1,\cr 0.25\pi,&\mbox{otherwise}.\end{array}\right.$
@@
We apply our positive scheme with adaptive viscosity (Algorithm 3) on irregular nodes obtained in the same way as in Example 2 by modifying $400^{2}$ Halton points in $[-2,2]\times[-2.5,1.5]$ near the boundaries. This corresponds to the spacing $h=0.01$, and hence we choose $\Delta t=0.2h=2\cdot 10^{-3}$, $\mu=0.5h=5\cdot 10^{-3}$. The results at $T=1$ are again compared with the solution obtained by employing the first order monotone scheme of Hesthaven [2018], Script 7.2, on the $400\times 400$ Cartesian grid. Figure 6.6(b,c) compares the solutions at the final time and the corresponding contour lines. The results reproduce the correct shape, with our solution having edges of discontinuity a bit less neat, which is expected having used Halton points. Despite this, Table 6.3 shows that the error with respect to the reference solution obtained with our method is somewhat better. Finally, Figure 6.7 shows the detected fault points at the final time, indicating that the detection algorithm behaves correctly.
@@
Figure 6.7: Example 3: Detected fault points (in red) at $T=1$ in the solution by the positive scheme with adaptive viscosity.
@@ 7 Conclusion
There are many directions of research of interest for future work. Purely polynomial numerical differentiation may be enhanced by kernel-based techniques that are used successfully in the meshless finite difference methods of RBF-FD type, see Fornberg and Flyer [2015], Davydov and Schaback [2016], Bayona et al. [2017], Davydov and Schaback [2019]. The flexibility of the space discretization makes it natural to explore in this setting the local time-stepping combined with local refinement in space. Positive schemes may be used locally where appropriate, enhanced by high order schemes as part of limiter- or ENO/WENO-based algorithms. In particular, an extension to the conservation laws of the meshless FCT method for linear convection dominated problems Sokolov et al. [2019] could be considered.
@@ Acknowledgments
C. Bracco, C. Giannelli and A. Sestini are members of the INdAM Research Group GNCS. This work has been partially supported by INdAM through Finanziamenti Premiali SUNRISE and GNCS.

### v2

@@ 6.2 Example 2: Inviscid Burgers’ equation with smooth initial condition
In spite of the initial condition being smooth, a shock appears as time advances. Note that this problem is an adaptation of the example given in Section 7.1.1 of Hesthaven [2018], where the flux is chosen slightly differently as ${\bf F}(u)=u^{2}{\bf v}$. While no explicit exact solution is available, we consider as a reference a high-resolution solution obtained by a fifth order WENO scheme on the $400\times 400$ Cartesian grid, see Chapter 11 of the book Hesthaven [2018] and its supplementary code BurgersWENO2D.m. For comparison with our method we choose a first order monotone scheme (Script 7.2 in Hesthaven [2018] and code BurgersM2D.m) on the $200\times 200$ Cartesian grid with spacing $h=0.0025$. Contour lines of the reference solution and the solution by the first order monotone method at time $T=0.1$ are shown in Figure 6.4(a,b).
@@
The contour lines of the solutions at time $T=0.1$ obtained by our method on the grid and Halton nodes are presented in Figure 6.4(c,d). In addition, we compare the cross sections of all five solutions for $x_{2}=0.1$ in Figure 6.5, and the errors with respect to the reference solution in Table 6.2. Note that in order to compute the cross sections and contour lines in the case of Halton or random nodes we interpolate the solution on the $200\times 200$ Cartesian grid by the piecewise linear polynomial on the Delaunay triangulation of the nodes. The results of the meshless scheme on the grid, Halton and random nodes are very similar, and significantly better than those of the monotone scheme. This is clear from the errors in the table and zoom-in plots in Figure 6.5(c-n). In particular, these results confirm remarkable robustness of our method with respect to the use of scattered points. Note that the contours in Figure 6.4(d) appear wobbling at the shocks, due to the irregular distribution of the Halton nodes in close proximity of the shock lines.
[…]
method $E_{1}$-error $E_{2}$-error monotone scheme (Hesthaven [2018], Script 7.2) 1.21e-02 4.38e-02 meshless scheme (Algorithm 3), Cartesian grid 6.08e-03 2.79e-02 meshless scheme (Algorithm 3), Halton points 5.47e-03 2.71e-02 meshless scheme (Algorithm 3), Random points 6.45e-03 3.34e-02
Table 6.2: Example 2: Comparison of the errors (with respect to the reference solution) obtained by a standard monotone scheme on the Cartesian grid, and our meshless method on the Cartesian grid, Halton and random nodes ($T=0.1$).
@@
(c) Grid: zoom no. 1.
(d) Grid: zoom no. 2.
(e) Grid: zoom no. 3.
(f) Grid: zoom no. 4.
(g) Halton: zoom no. 1.
(h) Halton: zoom no. 2.
(i) Halton: zoom no. 3.
(j) Halton: zoom no. 4.
(k) Random: zoom no. 1.
(l) Random: zoom no. 2.
(m) Random: zoom no. 3.
(n) Random: zoom no. 4.
Figure 6.5: Example 2: Comparison of the cross sections at $x_{2}=0.1$ of the reference solution and the monotone scheme with the meshless scheme of Algorithm 3 on (a) the Cartesian grid and (b) Halton points ($T=0.1$). Subfigures (c)–(n) show zoom-in plots of four critical portions of the cross sections near the shocks, comparing three meshless solutions (on the Cartesian grid, Halton and random points) to the reference solution and the solution by the monotone scheme.
(a) Spacing 0.01.
(b) Spacing 0.005.
(c) Spacing 0.0025.
(d) Spacing 0.00125.
Figure 6.6: Example 2: Comparison of the cross sections at $x_{2}=0.1$ of the reference solution and the monotone scheme with the meshless scheme of Algorithm 3 on Halton points ($T=0.1$) with decreasing spacing in the vicinity of the same corner as in the plots (e), (i) and (m) of Figure 6.5.
@@ 6.3 Example 3: Rotating wave
$u_{0}({\bf x})=\left\{\begin{array}[]{ll}3.5\pi,&\mbox{if }\|{\bf x}\|_{2}<1,\cr 0.25\pi,&\mbox{otherwise}.\end{array}\right.$
@@
We apply our positive scheme with adaptive viscosity (Algorithm 3) on irregular nodes obtained in the same way as in Example 2 by modifying $400^{2}$ Halton points in $[-2,2]\times[-2.5,1.5]$ near the boundaries. This corresponds to the spacing $h=0.01$, and hence we choose $\Delta t=0.2h=2\cdot 10^{-3}$, $\mu=0.5h=5\cdot 10^{-3}$. The results at $T=1$ are again compared with the solution obtained by employing the first order monotone scheme of Hesthaven [2018], Script 7.2, on the $400\times 400$ Cartesian grid. Figure 6.7(b,c) compares the solutions at the final time and the corresponding contour lines. The results reproduce the correct shape, with our solution having edges of discontinuity a bit less neat, which is expected having used Halton points. Despite this, Table 6.3 shows that the error with respect to the reference solution obtained with our method is somewhat better. Note that the maximum of the numerical solution is preserved at the same value $3.5\pi$ up to machine precision at all time steps. Finally, Figure 6.8 shows the detected fault points and the distribution of the viscosity values at the final time, indicating that viscosity is applied near the shocks as intended.
@@
(a) Fault points (in red).
(b) Values of $\mu_{i}$ obtained by formula (32).
Figure 6.8: Example 3: Detected fault points and viscosity values at $T=1$.
@@ 7 Conclusion
There are many directions of research of interest for future work. Purely polynomial numerical differentiation may be enhanced by kernel-based techniques that are used successfully in the meshless finite difference methods of RBF-FD type, see Fornberg and Flyer [2015], Davydov and Schaback [2016], Bayona et al. [2017], Davydov and Schaback [2019]. The flexibility of the space discretization makes it natural to explore in this setting the local time-stepping combined with local refinement in space. Positive schemes may be used locally where appropriate, enhanced by high order schemes as part of limiter- or ENO/WENO-based algorithms. In particular, an extension to the conservation laws of the meshless FCT method for linear convection dominated problems Sokolov et al. [2019] could be considered. An important goal for future work is the extension of the introduced positive schemes to general vector-valued conservation problems.
@@ Acknowledgments
We thank anonymous reviewers whose comments helped to improve the quality of the paper. C. Bracco, C. Giannelli and A. Sestini are members of the INdAM Research Group GNCS. This work has been partially supported by INdAM through Finanziamenti Premiali SUNRISE and GNCS.